The company has just hired a new marketing manager who insists that unit sales can be dramatically increased by dropping the selling price from $8 to $7. The marketing manager would like to use the following projections in the budget:

Year 2 Quarter

Year 3 Quarter

Data 1 2 3 4 1 2
Budgeted unit sales 45,000 65,000 105,000 60,000 90,000 100,000
Selling price per unit $7

8

9

10

11

12

13

14

15

16

17

18

19

Chapter 8: Applying Excel
Data Year 3 Quarter
1 2 3 4 1 2
Budgeted unit sales 45,000 65,000 105,000 60,000 90,000 100,000
• Selling price per unit $7 per unit
• Accounts receivable, beginning balance $65,000
• Sales collected in the quarter sales are made 75%
• Sales collected in the quarter after sales are made 25%
• Desired ending finished goods inventory is 30% of the budgeted unit sales of the next quarter
• Finished goods inventory, beginning 12,000 units
• Raw materials required to produce one unit 5 pounds
• Desired ending inventory of raw materials is 10% of the next quarter's production needs
• Raw materials inventory, beginning 23,000 pounds
• Raw material costs $0.80 per pound
• Raw materials purchases are paid 60% in the quarter the purchases are made
and 40% in the quarter following purchase
• Accounts payable for raw materials, beginning balance $81,500

1. What are the total expected cash collections for the year under this revised budget?

2. What is the total required production for the year under this revised budget?

3. What is the total cost of raw materials to be purchased for the year under this revised budget?

4. What are the total expected cash disbursements for raw materials for the year under this revised budget?
Answer #1

please find below table useful to compute desired results: -

A в с Q1 45000 102 Q2 65000 Year 285000 G Q1Yr-2 90000 H Q2Yr-2 100000 1 SALES BUDGET 2 Budgeted Sales units 3 Selling price

B Ε C 0.8 =C28*C29 08 =B28*B29 D 0.8 =D28*D29 10.8 0.8 =F28*F29 =E28 E29 02 А 29 Cost price per unit 30 Budgeted Purchase in

Ending results would appear as follows: -

1 A B C D E F G 1 SALES BUDGET Q1 Q2 Q3 Q4 Year Q1Yr-2 Q2Yr-2 2 Budgeted Sales units 45,000 65,000 105,000 60.000 285,000 90.

G B C D E F 0.8 0.8 0.8 0.8 0.8 222,400313,800 357.000 285,600 1,218,800|| А 29 Cost price per unit 30 Budgeted Purchase in $